OpenStudy (lgbasallote):

I don't get why this is zero..i got \(\infty\) \[\LARGE \int_{-\infty}^{\infty} xe^{-x^2} dx\]

OpenStudy (inkyvoyd):

Well, whats \[\int\limits_{}^{}xe ^{-x^2}dx\]

myininaya (myininaya):

Hint: \[\int\limits_{}^{}xe^{-x^2} dx \text{ Let } u=-x^2 => du=-2 x dx => \frac{-1}{2} du =dx\]

OpenStudy (lgbasallote):

hmm lemme type my solution..gonna be a lot of lateces...

OpenStudy (lgbasallote):

\[\lim_{t \rightarrow \infty} \int_{0}^{t} xe^{-x^2}dx + \lim_{t \rightarrow -\infty} \int_{-t}^{0} xe^{-x^2} dx\] \[\large -\frac{1}{2}\lim_{t \rightarrow \infty} [e^u] \stackrel{t}{0} - \frac{1}{2}\lim_{t \rightarrow -\infty} [e^u] \stackrel{0}{-t} \] by the way...it would be fun to learn how to do that latex other than \stackrel that looks like a fraction without the fraction line anyway \[ [e^{\infty} - e^0] = \infty\] \[e^0 - e^{-\infty}] = 1 - 0 = 1\] \[\frac{-\infty}{2} - \frac{1}{2} = -\infty\] right?

myininaya (myininaya):

\[\int\limits_{}^{}xe^{-x^2} dx=\int\limits_{}^{}e^u \frac{-1}{2} du=\frac{-1}{2}e^u+C=\frac{-1}{2}e^{-x^2}+C\] \[\int\limits_{-\infty}^{\infty} xe^{-x^2} dx=\int\limits_{-\infty}^{0}xe^{-x^2} dx+\int\limits_{0}^{\infty}xe^{-x^2} dx\] \[\lim_{a \rightarrow -\infty}[\frac{-1}{2}e^{-x^2}]_{a}^0+\lim_{b \rightarrow \infty}[\frac{-1}{2}e^{-x^2}]_0^{b}\] \[\lim_{a \rightarrow -\infty}[\frac{-1}{2}-\frac{-1}{2}e^{-a^2}]+\lim_{b \rightarrow \infty}[\frac{-1}{2}e^{-b^2}-\frac{-1}{2}]\] But \[\lim_{u \rightarrow \infty}e^{-u}=0\] So we have both parts converge :) \[\frac{-1}{2}-\frac{-1}{2}\]

myininaya (myininaya):

x^2 is positive no matter if x approaches negative infinity or positive infinity

myininaya (myininaya):

So -x^2 is always negative if x approaches negative infinity or positive infinity

myininaya (myininaya):

\[\lim_{u \rightarrow \pm \infty}e^{-u^2}=0\]

OpenStudy (lgbasallote):

uhmm okay...i think i stopped getting it after \[\lim_{a \rightarrow -\infty} [\frac{-1}{2} - \frac{1}{2}\] thingy what does that mean??

OpenStudy (lgbasallote):

uhmm @myininaya ?

OpenStudy (inkyvoyd):

Those are parenthesis

OpenStudy (inkyvoyd):

@myininaya just uses those to show that it's the limit of the whole thing, not just the part.

myininaya (myininaya):

Did that explain you question @Igbasallote because I didn't understand your question?

OpenStudy (lgbasallote):

what i dont get is where the second -1/2 sprouted from

myininaya (myininaya):

Plugin in the upper and lower limits?

OpenStudy (lgbasallote):

ohhhh thanks...ill go back to the solution again now

OpenStudy (lgbasallote):

i get it...seems the probelm with what i did is that i didnt bring back to -x^2 *facepalm*

myininaya (myininaya):

\[\frac{-1}{2}e^{-x^2}|_{a^2}^0=\frac{-1}{2}e^{-0^2}-\frac{-1}{2}e^{-a^2}=\frac{-1}{2}(1)+\frac{1}{2}e^{-a^2}=\frac{-1}{2}+\frac{1}{2}e^{-a^2}\]

myininaya (myininaya):

We do the other part similarly

OpenStudy (lgbasallote):

how do you do those limits o.O the one by the brackers

myininaya (myininaya):

Are you still talking about pluggin' in the upper and lower limit?

OpenStudy (lgbasallote):

no..the latex..how?

myininaya (myininaya):

You know you can look at what I type by clicking it and going to tex commands

OpenStudy (lgbasallote):

google chrome haha =)) it doesnt view there

myininaya (myininaya):

And it will show you what I typed to get the prettiness

myininaya (myininaya):

Like does it automatically minimize the window? And you can't open it?

myininaya (myininaya):

It really is annoying. I have chrome. I know what you are talking about but you can go to your chrome thingy at the task bar below (or to the side if you are that type) and if you maximize the screen you will see what I wrote OR I can just tell you lol

myininaya (myininaya):